# Data-Driven Linearization of Dynamical Systems  [^1] To appear in Nonlinear Dynamics

### v1

@@
Data-Driven Linearization of Dynamical Systems  [^1] To appear in Nonlinear Dynamics
@@ 1 Introduction
Despite these challenges, we may proceed to find an appropriately defined closest linear system (4) based on available observable data. We assume that for some fixed time step $\Delta t$, discrete observations of $m$ initial conditions, $\mathbf{x}^{1}(t_{0}),\ldots,\mathbf{x}^{m}(t_{0})$, and their images $\mathbf{F}^{t_{0}+\Delta t}(\mathbf{x}^{1}(t_{0})),\ldots,\mathbf{F}^{t_{0}+\Delta t}\left(\mathbf{x}^{m}(t_{0})\right)$, under the sampled flow map $\mathbf{F}^{t_{0}+\Delta t}$ are available in the data matrices
$\boldsymbol{\Phi}=\left[\boldsymbol{\phi}\left(\mathbf{x}^{1}(t_{0})\right),\ldots,\boldsymbol{\phi}\left(\mathbf{x}^{m}(t_{0})\right)\right],\qquad\hat{\boldsymbol{\Phi}}=\left[\boldsymbol{\phi}\left(\mathbf{F}^{t_{0}+\Delta t}(\mathbf{x}^{1}(t_{0}))\right),\ldots,\boldsymbol{\phi}\left(\mathbf{F}^{t_{0}+\Delta t}\left(\mathbf{x}^{m}(t_{0})\right)\right)\right],$ (5)
@@ 2 Prior justifications for DMD methods
Finally, a lesser known limitation of the Koopman-based approach to DMD is the limited domain in the phase space over which Koopman eigenfunctions (and hence their corresponding invariant subspaces) are defined in the observable space. Specifically, all $C^{\infty}$ Koopman eigenfunctions necessarily blow up near basin boundaries of attracting and repelling fixed points, periodic orbits and more general invariant sets (see Proposition 1 of our Appendix A.4 for a precise statement and proof).
Expansions of observables in terms of such blowing-up eigenfunctions have even smaller domains of convergence, as was shown explicitly in a simple example by Page and Kerswell [44]. This is a fundamental obstruction to the often envisioned concept of global linear models built of different Koopman eigenfunctions over multiple domains of attraction (see, e.g., Williams et al. [58], p. 1309). While it is broadly known that such models would be discontinuous along basin boundaries (Liu et al. [38, 39], Kvalheim and Arathoon [33]), it is rarely noted that such models would also generally blow up at those boundaries and hence would become unmanageable even before reaching the boundaries.
@@ 3.1 Justification of DMD for continuous dynamical systems
for the linearized dynamics, with linear decay rate towards $E$ strictly dominating all decay rates inside $E$. We also define the (real) spectral subspace of faster decaying linear modes:
@@
$\mathbf{X}=\left[\mathbf{x}^{1}(t_{0}),\ldots,\mathbf{x}^{m}(t_{0})\right],\quad\hat{\mathbf{X}}=\left[\mathbf{F}^{t_{0}+\Delta t}(\mathbf{x}^{1}(t_{0})),\ldots,\mathbf{F}^{t_{0}+\Delta t}\left(\mathbf{x}^{m}(t_{0})\right)\right]$ (18)
for trajectory data in the underlying dynamical system ((1)) on which the observable data matrices $\boldsymbol{\Phi}$ and $\hat{\boldsymbol{\Phi}}$ defined in (5) are defined. In truly data-driven applications, the matrices $\mathbf{X}$ and $\hat{\mathbf{X}}$ are not known. We will nevertheless use them to make precise statements about a required dominance of the slow linear modes of $E$ in the available data. Such a dominance will arise for generic initial conditions if one selects the initial conditions $\mathbf{x}^{1}(t_{0}),\ldots,\mathbf{x}^{m}(t_{0})$ after initial fast transients along $F$ have died out. This can be practically achieved by initializing $\mathbf{x}^{1}(t_{0}),\ldots,\mathbf{x}^{m}(t_{0})$ after a linear spectral analysis of the observable data matrix $\boldsymbol{\Phi}$ only returns only $d$ dominant frequencies.
@@
This means that the $\mathbf{x}=\mathbf{0}$ fixed point is only assumed hyperbolic with a spectral gap and $\mathbf{A}$ has an attracting $d$-dimensional spectral subspace $E$ that possibly contains some instabilities, i.e., eigenvalues with positive real parts. Then the statements of Theorem 1 still hold, but $\mathcal{W}(E)$ will be only be guaranteed $C^{1}$ at $\mathbf{x}=\mathbf{0}$ and Hölder-continuous at other points near the fixed point. This follows by replacing the linearization theorem of Hartman [23] with that of van Strien [56], which still enables us to use eq. (87) in the proof. Therefore, slow subspaces $E$ containing a mixture of stable and unstable modes can also be allowed, as long as $F$ contains only fast modes consistent with the splitting assume in eq. (12). In that case, however, the time $t_{0}+\Delta t$ must be chosen carefully to ensure that $\left|\mathbf{P}_{F}^{\mathrm{}}\hat{\mathbf{X}}\right|\leq\left|\mathbf{P}_{E}^{\mathrm{}}\mathbf{X}\right|^{1+\beta}$ still holds, i.e., the data used in DMD still samples a neighborhood of the origin.
@@ 4.1 Theoretical foundation for DDL
    $\mathbf{f}_{2}\in C^{r}$ and the nonresonance conditions
    $\lambda_{k}\neq\sum_{j=1}^{n}m_{j}\lambda_{j},\qquad j,k=1,\ldots,n,\qquad 2\leq\sum_{j=1}^{n}m_{j}\leq Q\leq r,\quad Q:=\mathrm{Int}\left[\frac{\max_{i}\left|\mathrm{Re}\,\lambda_{i}\right|}{\min_{i}\left|\mathrm{Re}\,\lambda_{i}\right|}\right]+1,$ (36)
    hold for the eigenvalues of $\mathbf{A}$.
@@
$\boldsymbol{\ell}(\boldsymbol{\gamma})=\sum_{\left|\mathbf{k}\right|=2}^{r}\mathbf{l}_{\mathbf{k}}\gamma^{\mathbf{k}}+o\left(\left|\boldsymbol{\gamma}\right|^{r}\right),\quad\mathbf{k}\in\mathbb{N}^{d},\quad\mathbf{l}_{\mathbf{k}}\in\mathbb{R}^{d},\quad\boldsymbol{\gamma}^{\mathbf{k}}:=\gamma_{1}^{k_{1}}\cdots\gamma_{d}^{k_{d}}.$ (42)
@@
$\boldsymbol{\ell}(\boldsymbol{\gamma})=\sum_{\left|\mathbf{k}\right|=2}^{\infty}\mathbf{l}_{\mathbf{k}}\gamma^{\mathbf{k}},\quad\mathbf{k}\in\mathbb{N}^{d},\quad\mathbf{l}_{\mathbf{k}}\in\mathbb{R}^{d},\quad\boldsymbol{\gamma}^{\mathbf{k}}:=\gamma_{1}^{k_{1}}\cdots\gamma_{d}^{k_{d}}$ (43)
@@
Assuming hyperbolicity is essential for Theorem ‘4 to hold, since in this case the linearization is the same as transforming the dynamics to the Poincaré-normal form. For a non-hyperbolic fixed point, this normal form transformation results in nonlinear dynamics on the center manifold. This would, however, only arise in highly non-generic systems, precisely tuned to be at criticality. Since this is unlikely to happen in experimentally observed or numerically simulated systems, the hyperbolicity assumption is not restrictive.
@@ 5 Examples
For $R\geq 0$, the system has a repelling fixed point at $R=0$ and an attracting one at $R=1$. Page and Kerswell [44] show that local expansion of observables in terms of the Koopman eigenfunctions computed near each fixed point are possible, but the expansions at the two fixed points are not compatible with each other and both diverge at $R=\sqrt{2}/2$. This is a consequence of the more general result that the Koopman eigenfunctions themselves inevitably blow up near basin boundaries (see our Proposition 1 in Appendix A of the Supplementary Information). Both DMD and EDMD can nevertheless be computed from data, even for a trajectory crossing the turning point at $R=\sqrt{2}/2$, but the resulting models cannot have any connection to the Koopman operator.
In each comparison performed on system (66), we generate a single trajectory in the domain of attraction of the $R=1$ fixed point and use it as training data for DMD, EDMD and DDL. In each subplot of Fig. 3, the single training trajectory starts from the intersection of the red horizontal line “IC of training trajectory” with the black curve labeled “True trajectory” at time $t=0$. We then also generate a new test trajectory (black) with its initial condition denoted with a black dot over the line $t=0$. We place this initial condition slightly outside the domain of linearization for system (66) (under the grey line labelled “Turning point”). We use DMD, order-$k=5$ EDMD, and DDL trained on a single training trajectory to make predictions for the black testing trajectory (not used in the training).
@@
The predictions of DMD and EDMD can only be considered accurate for very low amplitude oscillations, while DDL returns accurate predictions throughout the whole trajectory. This example consists of linear dynamics and monomial observables of the state, and hence should be an ideal test case for EDMD. Yet, EDMD is inaccurate in identifying the spectrum of system (68). Indeed, as seen in Fig. 4b, a number of spurious eigenvalues arise from EDMD, both real and complex, and only one can be considered to be close to a true eigenvalue. DMD performs clearly better but it is still markedly less accurate than DDL. These inaccuracies in the predictions of EDMD and DMD spectra are also reflected by considerable errors in their predictions for trajectories, as seen in Fig. 4a. In contrast, DDL produces the most accurate prediction for the test trajectory.
@@ 5.4 Water sloshing experiment in a tank
All this analysis has been carried out using the publicly available SSMLearn package Cenedese et al. [11]. With $\mathcal{W}\left(E\right)$ identified, we use the DDL method with order $k=5$ to find the linearizing transformation and the linearized dynamics on $\mathcal{W}\left(E\right)$. In Fig 7(b) we show the prediction of the model on a decaying trajectory reserved for testing. The displacements are reported as percentage values, with respect to the depth of the tank. In Fig. 7c-d, we show predictions from the DDL model for the forced response, compared with the experimentally observed response. Since the exact forcing function is unknown, we follow the calibration procedure outlined by Cenedese et al. [12] to find an equivalent forcing amplitude in the reduced-order model.
We present data for three forcing amplitudes. The DDL predictions are accurate up to $0.17\%$ amplitude, even capturing the softening trend. The largest-amplitude forcing resulted in response significantly outside the range of the training data; in this range, we were unable to find the converged forced response from DDL. We also show the corresponding DMD-predictions In Fig. 7c-d. Although the linear response can formally be evaluated for any forcing amplitude, DMD shows no trace of the softening trend, and is even inaccurate for low forcing amplitudes.
Figure 7: (a) Schematic representation of the experimental setup (adopted from Cenedese et al. [12]). (b) Prediction of the decay of a test trajectory with order-$k=5$ DDL. (c) Prediction of the forced response from DMD. (d) Prediction of the forced response from DDL. Blue shading indicates the domain, in which training data for DMD and DDL was available.
@@ 5.5 Model reduction and foliation in a nonlinear oscillator chain
Three numerically generated training trajectories show decay to the $\mathbf{q}=\mathbf{0}$ fixed point, as expected from the damped nature of the linear part of the system. In this example, we also seek to capture some of the transients, which motivates us to select the slow SSM $\mathcal{W}(E)$ to be 4D, tangent to the spectral subspace $E=E_{1}\oplus E_{2}$ spanned by the the slowest mode ($E_{1})$ and the second slowest mode ($E_{2}$). As the mode corresponding to $E_{2}$ does disappear over time from the decaying signal, there is no resonance between the eigenvalues and hence Theorem 4 is applicable. As already noted, numerical from a generic physical system described by eq. (74) will be free from resonances. An exception is a $1\colon 1$ resonance arising from a perfect symmetry, but this resonance is not excluded by Theorem 4 and has is amenable to DDL.
@@ 6 Conclusions
Data Availability All data and codes used in this work will be made available after publication.
@@
Leonhardstrasse 21, 8092 Zürich, Switzerland
@@ A.4 Koopman eigenfunctions
All $C^{\infty}$ Koopman eigenfunctions blow up (i.e., become unbounded) at the boundaries of domains of attraction or repulsion of a compact invariant set along which the underlying dynamical system admits a $C^{1}$linearization transformation.
@@
We only prove the statement for domains of attraction, as domains of repulsion can be handled in the same fashion in backward time. Let $\mathbf{x}_{0}$ be a point on the boundary $\mathcal{B}$ of a domain of attraction of a compact invariant manifold $\mathcal{S}$. Examples of such include attracting fixed points, periodic orbits and quasiperiodic tori. As the domain of attraction is open, for any $\rho>0$, there is a closed $\rho$-neighborhood $B_{\rho}(\mathcal{S})$ around the invariant set $\mathcal{S}$ that lies fully in the domain of attraction. Consequently, we have $d(\mathcal{S},\mathbf{F}^{t}(\mathbf{x}_{0}))>\rho/2>0$ for all $t\geq 0$. By continuous dependence of the flow map on initial conditions, for any $T>0$ there exists $\delta(T)>0$ such that for all $\hat{\mathbf{x}}_{0}\in B_{\delta(T)}\left(\mathbf{x}_{0}\right)\cap\mathcal{B}$, we have
$d(\mathbf{F}^{t}(\hat{\mathbf{x}}_{0}),\mathbf{F}^{t}(\mathbf{x}_{0}))<\rho/2,\quad t\in\left[0,T\right].$
This means that for arbitrary large $T$, no initial condition in $B_{\delta(T)}\left(\mathbf{x}_{0}\right)\cap\mathcal{B}$ will reach the ball $B_{\rho}(\mathcal{S})$ around the origin for times up to $T$. This in turn means that for any $\mathbf{x}_{0}\in\mathcal{B},$ there are arbitrarily close initial conditions $\hat{\mathbf{x}}_{0}$ in the domain of attraction with arbitrarily long time of flight to the ball $B_{\rho}(\mathcal{S})$ around the attracting invariant set $\mathcal{S}.$
For the special case of a fixed point $p=\mathcal{S}$, Lan and Mezic [35] extended the local $C^{1}$ linearization result of Hartman [23] near an attracting fixed point via the backward-time flow map to a global $C^{1}$ linearizing transformation $\mathbf{x}=\mathbf{h}(\mathbf{y})$ within $\mathcal{B}$ that maps the original ODE within $\mathcal{B}$ to $\dot{\mathbf{y}}=\mathbf{A}\mathbf{y}$ defined on $\mathbb{R}^{n}$. This transformation has also been shown unique by Eldering et al. [17]. By the construction of this unique linearizing transformation, initial conditions $\hat{\mathbf{x}}_{0}$ whose times of flight to $B_{\rho}(p)$ is large will then map those initial conditions $\hat{\mathbf{y}}_{0}$ of the linearized system that are far from the origin $\mathbf{y}=\mathbf{0}$.
@@
$\boldsymbol{\phi}_{j}(\mathbf{x})=\left\langle\mathbf{h}^{-1}(\mathbf{x}),\mathbf{v}_{j}\right\rangle,\quad j=1,\ldots,n.$ (86)
As we have just concluded that $\mathbf{y}=\mathbf{h}^{-1}(\mathbf{x})$ will admit arbitrarily large values arbitrarily close to any point $\mathbf{x}_{0}\in\mathcal{B},$ the principal Koopman eigenfunctions $\boldsymbol{\phi}_{j}(\mathbf{x})$ will also have arbitrarily large values near any $\mathbf{x}_{0}\in\mathcal{B}$ by eq. (86). (We need to exclude points $\hat{\mathbf{x}}_{0}$ at which $\mathbf{h}^{-1}(\hat{\mathbf{x}_{0}})\perp\mathbf{v}_{j}$, but those points form an $n-1$-dimensional plane $\mathcal{V}(\hat{\mathbf{x}}_{0})$ that has a measure zero intersection with $B_{\delta(T)}\left(\mathbf{x}_{0}\right)\cap\mathcal{B}$.) Therefore, the set $B_{\delta(T)}\left(\mathbf{x}_{0}\right)-\mathcal{V}(\hat{\mathbf{x}}_{0})$ still has arbitrarily many points with arbitrarily large values of $\boldsymbol{\phi}_{j}(\mathbf{x})$. This establishes the blow-up of any principal Koopman eigenfunction $\boldsymbol{\phi}_{j}(\mathbf{x})$ at any of the domain boundary points $\mathbf{x}_{0}\in\mathcal{B}$.
Kvalheim and Revzen [34] show that all $C^{\infty}$ Koopman eigenfunctions are products of powers of principal Koopman eigenfunctions. As a consequence, we obtain that all $C^{\infty}$ Koopman eigenfunctions blow up a the boundary $\mathcal{B}$ of the domain of attraction of a fixed point. Finally, Eldering et al. [17] extend the results of Lan and Mezic [35] from attracting fixed points $p$ to arbitrary compact invariant manifolds $\mathcal{S}$ with a domain of attraction, as long as a local $C^{1}$ linearizing transformation exists near $\mathcal{S}.$

### v2

@@
Data-Driven Linearization of Dynamical Systems
@@ 1 Introduction
Despite these challenges, we may proceed to find an appropriately defined closest linear system (4) based on available observable data. We assume that for some fixed time step $\Delta t$, discrete observations of $m$ initial conditions, $\mathbf{x}^{1}(t_{0}),\ldots,\mathbf{x}^{m}(t_{0})$, and their images $\mathbf{F}^{\Delta t}(\mathbf{x}^{1}(t_{0})),\ldots,\mathbf{F}^{\Delta t}\left(\mathbf{x}^{m}(t_{0})\right)$, under the sampled flow map $\mathbf{F}^{\Delta t}$ are available in the data matrices
$\boldsymbol{\Phi}=\left[\boldsymbol{\phi}\left(\mathbf{x}^{1}(t_{0})\right),\ldots,\boldsymbol{\phi}\left(\mathbf{x}^{m}(t_{0})\right)\right],\qquad\hat{\boldsymbol{\Phi}}=\left[\boldsymbol{\phi}\left(\mathbf{F}^{\Delta t}(\mathbf{x}^{1}(t_{0}))\right),\ldots,\boldsymbol{\phi}\left(\mathbf{F}^{\Delta t}\left(\mathbf{x}^{m}(t_{0})\right)\right)\right],$ (5)
@@ 2 Prior justifications for DMD methods
Finally, a lesser known limitation of the Koopman-based approach to DMD is the limited domain in the phase space over which Koopman eigenfunctions (and hence their corresponding invariant subspaces) are typically defined in the observable space. Specifically, at least one principal Koopman eigenfunction necessarily blows up near basin boundaries of attracting and repelling fixed points and periodic orbits (see Proposition 1 of our Appendix A.4 for a precise statement and Theorem 3 of Kvalheim and Arathoon [33] for a more general related result).
Expansions of observables in terms of such blowing-up eigenfunctions have even smaller domains of convergence, as was shown explicitly in a simple example by Page and Kerswell [44]. This is a fundamental obstruction to the often envisioned concept of global linear models built of different Koopman eigenfunctions over multiple domains of attraction (see, e.g., Williams et al. [58], p. 1309). While it is broadly known that such models would be discontinuous along basin boundaries (Liu et al. [38, 39], Kvalheim and Arathoon [33]), it is rarely noted (see Kvalheim and Arathoon [33] for such a rare exception) that such models would also generally blow up at those boundaries and hence would become unmanageable even before reaching the boundaries.
@@ 3.1 Justification of DMD for continuous dynamical systems
for the linearized dynamics, with linear decay rate towards $E$ strictly dominating all decay rates inside $E$. Note that the set of vectors ${\mathrm{Re}\ \mathbf{e}_{1},\mathrm{Im}\ \mathbf{e}_{1},…,\mathrm{Re}\ \mathbf{e}_{d},\mathrm{Im}\ \mathbf{e}_{d}}$ is, in general, not linearly independent, but they span a $d$-dimensional subspace. We also define the (real) spectral subspace of faster decaying linear modes:
@@
$\mathbf{X}=\left[\mathbf{x}^{1}(t_{0}),\ldots,\mathbf{x}^{m}(t_{0})\right],\quad\hat{\mathbf{X}}=\left[\mathbf{F}^{\Delta t}(\mathbf{x}^{1}(t_{0})),\ldots,\mathbf{F}^{\Delta t}\left(\mathbf{x}^{m}(t_{0})\right)\right]$ (18)
for trajectory data in the underlying dynamical system ((1)) on which the observable data matrices $\boldsymbol{\Phi}$ and $\hat{\boldsymbol{\Phi}}$ defined in (5) are defined. In truly data-driven applications, the matrices $\mathbf{X}$ and $\hat{\mathbf{X}}$ are not known. We will nevertheless use them to make precise statements about a required dominance of the slow linear modes of $E$ in the available data. Such a dominance will arise for generic initial conditions if one selects the initial conditions $\mathbf{x}^{1}(t_{0}),\ldots,\mathbf{x}^{m}(t_{0})$ after initial fast transients along $F$ have died out. This can be practically achieved by initializing $\mathbf{x}^{1}(t_{0}),\ldots,\mathbf{x}^{m}(t_{0})$ after a linear spectral analysis of the observable data matrix $\boldsymbol{\Phi}$ only returns a number of dominant frequencies consistent with a $d$-dimensional SSM.
@@
This means that the $\mathbf{x}=\mathbf{0}$ fixed point is only assumed hyperbolic with a spectral gap and $\mathbf{A}$ has an attracting $d$-dimensional spectral subspace $E$ that possibly contains some instabilities, i.e., eigenvalues with positive real parts. Then the statements of Theorem 1 still hold, but $\mathcal{W}(E)$ will only be guaranteed $C^{1}$ at $\mathbf{x}=\mathbf{0}$ and Hölder-continuous at other points near the fixed point. This follows by replacing the linearization theorem of Hartman [23] with that of van Strien [56], which still enables us to use eq. (87) in the proof. Therefore, slow subspaces $E$ containing a mixture of stable and unstable modes can also be allowed, as long as $F$ contains only fast modes consistent with the splitting assume in eq. (12). In that case, however, the time $t_{0}+\Delta t$ must be chosen carefully to ensure that $\left|\mathbf{P}_{F}^{\mathrm{}}\hat{\mathbf{X}}\right|\leq\left|\mathbf{P}_{E}^{\mathrm{}}\mathbf{X}\right|^{1+\beta}$ still holds, i.e., the data used in DMD still samples a neighborhood of the origin.
@@ 4.1 Theoretical foundation for DDL
    $\mathbf{f}_{2}\in C^{r}$ and for all nonnegative integers $m_{j}$, the nonresonance conditions
    $\lambda_{k}\neq\sum_{j=1}^{n}m_{j}\lambda_{j},\qquad k=1,\ldots,n,\qquad 2\leq\sum_{j=1}^{n}m_{j}\leq Q\leq r,\quad Q:=\left\lfloor\frac{\max_{i}\left|\mathrm{Re}\,\lambda_{i}\right|}{\min_{i}\left|\mathrm{Re}\,\lambda_{i}\right|}\right\rfloor+1,$ (36)
[…]
@@
$\boldsymbol{\ell}(\boldsymbol{\gamma})=\sum_{\left|\mathbf{k}\right|=2}^{r}\mathbf{l}_{\mathbf{k}}\boldsymbol{\gamma}^{\mathbf{k}}+o\left(\left|\boldsymbol{\gamma}\right|^{r}\right),\quad\mathbf{k}\in\mathbb{N}^{d},\quad\mathbf{l}_{\mathbf{k}}\in\mathbb{R}^{d},\quad\boldsymbol{\gamma}^{\mathbf{k}}:=\gamma_{1}^{k_{1}}\cdots\gamma_{d}^{k_{d}}.$ (42)
@@
$\boldsymbol{\ell}(\boldsymbol{\gamma})=\sum_{\left|\mathbf{k}\right|=2}^{\infty}\mathbf{l}_{\mathbf{k}}\boldsymbol{\gamma}^{\mathbf{k}},\quad\mathbf{k}\in\mathbb{N}^{d},\quad\mathbf{l}_{\mathbf{k}}\in\mathbb{R}^{d},\quad\boldsymbol{\gamma}^{\mathbf{k}}:=\gamma_{1}^{k_{1}}\cdots\gamma_{d}^{k_{d}}$ (43)
@@
Assuming hyperbolicity is essential for Theorem 4 to hold, since in this case the linearization is the same as transforming the dynamics to the Poincaré-normal form. For a non-hyperbolic fixed point, this normal form transformation results in nonlinear dynamics on the center manifold. This would, however, only arise in highly non-generic systems, precisely tuned to be at criticality. Since this is unlikely to happen in experimentally observed or numerically simulated systems, the hyperbolicity assumption is not restrictive.
@@ 5 Examples
For $R\geq 0$, the system has a repelling fixed point at $R=0$ and an attracting one at $R=1$. Page and Kerswell [44] show that local expansion of observables in terms of the Koopman eigenfunctions computed near each fixed point are possible, but the expansions at the two fixed points are not compatible with each other and both diverge at $R=\sqrt{2}/2$. This is a consequence of the more general result that the Koopman eigenfunctions themselves typically blow up near basin boundaries (see our Proposition 1 in Appendix A of the Supplementary Information). Both DMD and EDMD can nevertheless be computed from data, even for a trajectory crossing the turning point at $R=\sqrt{2}/2$, but the resulting models cannot have any connection to the Koopman operator.
In each comparison performed on system (66), we generate a single trajectory in the domain of attraction of the $R=1$ fixed point and use it as training data for DMD, EDMD and DDL. In each subplot of Fig. 3, the single training trajectory starts from the intersection of the red horizontal line “IC of training trajectory” with the $t=0$ dashed line. We then also generate a new test trajectory (black) with its initial condition denoted with a black dot over the line $t=0$. We place this initial condition slightly outside the domain of linearization for system (66) (under the grey line labeled “Turning point”). We use DMD, order-$k=5$ EDMD, and DDL trained on a single training trajectory to make predictions for the black testing trajectory (not used in the training).
@@
The predictions of DMD and EDMD can only be considered accurate for very low amplitude oscillations, while DDL returns accurate predictions throughout the whole trajectory. This example consists of linear dynamics and monomial observables of the state, and hence should be an ideal test case for EDMD. Yet, EDMD is inaccurate in identifying the spectrum of system (68). Indeed, as seen in Fig. 4b, a number of spurious eigenvalues arise from EDMD, both real and complex. DMD performs clearly better but it is still markedly less accurate than DDL. These inaccuracies in the predictions of EDMD and DMD spectra are also reflected by considerable errors in their predictions for trajectories, as seen in Fig. 4a. In contrast, DDL produces the most accurate prediction for the test trajectory.
@@ 5.4 Water sloshing experiment in a tank
All this analysis has been carried out using the publicly available SSMLearn package Cenedese et al. [11]. With $\mathcal{W}\left(E\right)$ identified, we use the DDL method with order $k=5$ to find the linearizing transformation and the linearized dynamics on $\mathcal{W}\left(E\right)$. In Fig 7(b) we show the prediction of the model on a decaying trajectory reserved for testing. The displacements are reported as percentage values, with respect to the depth of the tank. In Fig. 7c-d, we show predictions from the DMD and DDL models for the forced response, compared with the experimentally observed response. Since the exact forcing function is unknown, we follow the calibration procedure outlined by Cenedese et al. [12] to find an equivalent forcing amplitude in the reduced-order model.
We present data for three forcing amplitudes. The DDL predictions are accurate up to $0.17\%$ amplitude, even capturing the softening trend. The largest-amplitude forcing resulted in response significantly outside the range of the training data; in this range, we were unable to find the converged forced response from DDL. We also show the corresponding DMD-predictions in Fig. 7c. Although the linear response can formally be evaluated for any forcing amplitude, DMD shows no trace of the softening trend, and is even inaccurate for low forcing amplitudes.
Figure 7: (a) Schematic representation of the experimental setup (adopted from Cenedese et al. [12]). (b) Prediction of the decay of a test trajectory with order-$k=5$ DDL. (c) Prediction of the forced response from DMD. (d) Prediction of the forced response from DDL. Light shading indicates the domain, in which training data for DMD and DDL was available.
@@ 5.5 Model reduction and foliation in a nonlinear oscillator chain
Three numerically generated training trajectories show decay to the $\mathbf{q}=\mathbf{0}$ fixed point, as expected from the damped nature of the linear part of the system. In this example, we also seek to capture some of the transients, which motivates us to select the slow SSM $\mathcal{W}(E)$ to be 4D, tangent to the spectral subspace $E=E_{1}\oplus E_{2}$ spanned by the the slowest mode ($E_{1})$ and the second slowest mode ($E_{2}$). As the mode corresponding to $E_{2}$ does disappear over time from the decaying signal, there is no resonance between the eigenvalues and hence Theorem 4 is applicable. As already noted, numerical data from a generic physical system described by eq. (74) will be free from resonances. An exception is a $1\colon 1$ resonance arising from a perfect symmetry, but this resonance is not excluded by Theorem 4 and has is amenable to DDL.
@@ 6 Conclusions
Acknowledgement We are grateful to Matthew Kvalheim and Shai Revzen for several helpful comments on an earlier version of this manuscript.
@@
Data Availability All data and codes used in this work are downloadable from the repository https://github.com/haller-group/DataDrivenLinearization.
@@
Leonhardstrasse 76, 8092 Zürich, Switzerland
@@ A.4 Koopman eigenfunctions
At least one principal Koopman eigenfunction blows up (i.e., becomes unbounded) at the boundary of a domain of attraction or repulsion of a fixed point around which the underlying dynamical system admits a local $C^{1}$linearization.
@@
We only prove the statement for domains of attraction of fixed points of continuous dynamical systems. Domains of repulsion can be handled in the same fashion in backward time and the proof for discrete dynamical systems is similar. Let $\mathbf{x}_{0}$ be a point on the boundary $\partial\mathcal{B}$ of a domain of attraction $\mathcal{B}$ of a fixed point $p$ of a continuous dynamical system. Then, arbitrarily close to $\mathbf{x}_{0}$, there are initial conditions $\hat{\mathbf{x}}_{0}$ in the domain of attraction with arbitrarily long times of flight to the ball $B_{\rho}(p)$ around $p$ in which the dynamical system can be $C^{1}$linearized.
Lan and Mezic [35] extended local $C^{1}$ linearization near an attracting fixed point via the backward-time flow map to a global $C^{1}$ linearizing transformation $\mathbf{x}=\mathbf{h}(\mathbf{y})$ within $\mathcal{B}$ that maps the original ODE within $\mathcal{B}$ to $\dot{\mathbf{y}}=\mathbf{A}\mathbf{y}$ defined on $\mathbb{R}^{n}$. By the construction of this linearizing transformation, initial conditions $\hat{\mathbf{x}}_{0}$ whose times of flight to $B_{\rho}(p)$ is large will be mapped into initial conditions $\hat{\mathbf{y}}_{0}$ of the linearized system that are far from the origin $\mathbf{y}=\mathbf{0}$.
@@
$\boldsymbol{\phi}_{j}(\mathbf{x})=\left\langle\mathbf{h}^{-1}(\mathbf{x}),\mathbf{v}_{j}\right\rangle,\quad j=1,\ldots,n,$ (86)
with $\mathbf{v}_{j}$ denoting the left eigenvectors or $\mathbf{A}$. As we have just concluded that $\mathbf{y}=\mathbf{h}^{-1}(\mathbf{x})$ will admit arbitrarily large values arbitrarily close to any point $\mathbf{x}_{0}\in\mathcal{\partial\mathcal{B}}$, it follows that there will be at least one $j$ for which the principal Koopman eigenfunction $\boldsymbol{\phi}_{j}(\mathbf{x})$ blows up at $\mathbf{x}_{0}\in\partial\mathcal{B}$. ∎
We note that the statement of Proposition 1 can also be deduced from the more general results in Theorem 3 of Kvalheim and Arathoon [33].